Required information

The Foundational 15 [LO10-1, LO10-2]

[The following information applies to the questions displayed below.]

Westerville Company reported the following results from last year’s operations:

Sales $ 1,000,000
Variable expenses 300,000
Contribution margin 700,000
Fixed expenses 500,000
Net operating income $ 200,000
Average operating assets $ 625,000


At the beginning of this year, the company has a $120,000 investment opportunity with the following cost and revenue characteristics:

Sales $ 200,000
Contribution margin ratio 60 % of sales
Fixed expenses $ 90,000

The company’s minimum required rate of return is 15%.

Foundational 10-9

9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this year? (Do not round intermediate calculations. Round your percentage answer to 1 decimal place (i.e. 0.1234 should be entered as 12.3).)
Answer #1
Required 9 :
This year's investment opportunity:
Sales 200000
(-) Variable expenses (40%) 80000
Contribution margin (60%) 120000
(-) Fixed expenses 90000
Net operating income 30000
Total average operating assets = 625000 + 120000 745000
Total net operating income = 200000 + 30000 230000
ROI of this year = Total net operating income / Total average operating assets = 230000 / 745000    30.9%